OpenStudy (anonymous):

solve |9x+2| = |x-5|

jimthompson5910 (jim_thompson5910):

if |9x+2| = |x-5| then 9x+2 = x-5 or 9x+2 = -(x-5) solve each for x

OpenStudy (anonymous):

i have tried i cant get it

jimthompson5910 (jim_thompson5910):

ok I'll solve 9x+2 = x-5 for you and hopefully that's enough to get you started 9x+2 = x-5 9x = x-5-2 9x-x = -5-2 8x = -7 x = -7/8 So one solution is x = -7/8

jimthompson5910 (jim_thompson5910):

I'll let you solve the second equation

OpenStudy (anonymous):

would it be 3/10?

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

the two solutions are 3/10, -7/8

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome